### v1

# Curriculum-Guided Layer Scaling for Language Model Pretraining

Abstract
As the cost of pretraining large language models grows, there is continued interest in strategies to improve learning efficiency during this core training stage. Motivated by cognitive development, where humans gradually build knowledge as their brains mature, we propose Curriculum-Guided Layer Scaling (CGLS), a framework for compute-efficient pretraining that synchronizes increasing data difficulty with model growth through progressive layer stacking (i.e. gradually adding layers during training). At the 100M parameter scale, using a curriculum transitioning from synthetic short stories to general web data, CGLS outperforms baseline methods on the question-answering benchmarks PIQA and ARC. Pretraining at the 1.2B scale, we stratify the DataComp-LM corpus with a DistilBERT-based classifier and progress from general text to highly technical or specialized content. Our results show that progressively increasing model depth alongside sample difficulty leads to better generalization and zero-shot performance on various downstream benchmarks. Altogether, our findings demonstrate that CGLS unlocks the potential of progressive stacking, offering a simple yet effective strategy for improving generalization on knowledge-intensive and reasoning tasks.
1 Introduction
Large language models (LLMs) are typically pretrained in a single, continuous pass, processing all tokens with a uniform amount of computation regardless of their complexity or relevance to downstream tasks of interest. While this approach has shown remarkable success in large-scale models like GPT-4 [34] and Llama 3 [9], it differs significantly from how humans learn, often leading to models that excel in generating coherent text but struggle with long-context reasoning across varied tasks [40]. Recent works like Phi-3 [1], MiniCPM [24], and others [14] have explored midtraining, adjusting the training data distribution partway through training by incorporating higher-quality, multilingual, or long-form text. However, this coarse-grained curriculum is applied on fixed model architectures. Inspired by how humans progressively build knowledge alongside their physically growing brains, we explore whether gradually scaling a model in tandem with increasingly complex data can enable more efficient and effective learning.
Curriculum learning [4] has shown success in guiding models from easier to harder tasks, but remains largely unexplored in modern language model pretraining. Recent studies have proposed several alternative methods to schedule complexity and model scale during transformer training. Fan and Jaggi [12] train small proxy models to determine sample learnability, enabling efficient complexity-aware curricula for training large LMs. Progressive layer stacking [16, 25, 39, 18, 48, 35], layer-dropping [51], and pruning [28] are gradual architectural adjustments that accelerate training and reduce inference cost. In computer vision, progressive model growing has also been explored for image generation [27] and segmentation [15].
In particular, progressive stacking is an approach that trains a model in stages, adding new layers (randomly initialized or copied) at the start of each stage until a final model size is eventually reached. Despite the intuitive appeal of progressive stacking as an approach to learn hierarchical features and improve compute efficiency, recent studies suggest it underperforms in knowledge-intensive tasks such as closed-book question answering. For example, MIDAS [39] finds that progressively grown models lag behind full-capacity models trained from scratch on memorization-focused tasks. In our own setup and compute-controlled experiments, we also find that progressive growing performs similarly or worse than standard training on several benchmarks.
We hypothesize that model expansion alone is insufficient for compute-efficient learning of hierarchical features, and investigate whether stacking should be paired with a structured learning signal to realize its potential. Inspired by cognitive development, we propose Curriculum-Guided Layer Scaling (CGLS), a pretraining paradigm that couples progressive model expansion with a data curriculum of gradually increasing complexity. To preserve and build on previously learned representations during model growth, CGLS uses staged training: at each step, newly added transformer layers are first trained in isolation (with earlier layers frozen), before the entire model is fine-tuned on a more complex data distribution. This strategy protects previously learned representations during capacity increases and aligns model complexity with data difficulty throughout training.
Our Contributions.
In summary, we make the following contributions:
  • We propose Curriculum-Guided Layer Scaling (CGLS), a unified framework for scaling both model capacity and data complexity during pretraining, combining layerwise expansion and stage-wise transfer with curriculum learning.
  • We conduct a rigorous empirical evaluation of CGLS at two parameter counts and compute scales, demonstrating its scalability.
    • At a GPT-2-Small parameter count using a synthetic-to-webtext curriculum, CGLS achieves consistent gains in perplexity and downstream QA performance (1.0% on average, and +2.1% on PIQA).
    • At a LLaMA-3.2-1B parameter count, we stratify the DataComp-LM (DCLM) corpus [33] using a DistilBERT-based classifier [38] and show that CGLS outperforms compute-matched baselines on knowledge-intensive and reasoning-heavy tasks, on average by 2.17% across all benchmarks and by as much as 5% on ARC-Easy.
Overall, we demonstrate that joint scaling of model and data complexity unlocks the benefits of progressive stacking for language model pretraining, yielding better downstream performance on knowledge-intensive tasks up to the 1B parameter scale.
Related Work.
Curriculum learning, formalized by Bengio et al. [4], has shown empirical benefits in many domains, from vision [21, 45] to reinforcement learning [17]. However, it remains underexplored in the context of large language model (LLM) pretraining, where most models are trained in a single pass over a diverse and unstructured dataset. Self-paced learning strategies [31] adjust training dynamics based on model uncertainty, gradually introducing harder samples as the model becomes more confident. More recently, efforts such as DataComp [33], MetaCLIP [47], and Tower [2] demonstrate that large performance gains can arise from careful dataset curation, even without architectural changes. Related work like Thrush et al. [43] and Ankner et al. [3] investigates the use of small language models to identify high-quality data subsets that better align with LLM performance, but does not explore staged training. Likewise, Gururangan et al. [20] show the value of domain-specific continued pretraining after generic training, indicating that structured data exposure plays a key role even in later stages. Still, these approaches do not consider whether data curricula coordinated with model growing could improve compute efficiency.
Parallel to this, a growing body of research examines how to progressively increase model capacity during training. Soviany et al. [42] and others have noted that the curriculum principle may apply not just to data, but also to model size. In vision, progressive training has been used to stabilize training and reduce compute costs, e.g., by gradually growing generative networks [27] or using function-preserving transformations [7]. For Transformers, techniques such as LayerDrop [11] and Width-Adaptive Transformers [52] enable dynamic capacity at inference time, though they are not designed for staged training. Mixture-of-Experts (MoE) models [41, 13] similarly allocate compute conditionally across tokens but often suffer from expert collapse or shallow specialization to token-level features [54]. More recently, architectural methods have emerged to progressively grow models during training. Wang et al. [44] propose learning an explicit mapping from a smaller transformer’s parameters into a larger one, enabling weight reuse and transfer. Complementary to this, and building on the previously-cited literature on gradual stacking, Yao et al. [49] introduce smoother transitions between training stages by gradually increasing the influence of new parameters. Such efforts underscore the promise of synchronized curricula for data and model: gradually increasing data complexity or domain diversity in tandem with growing model capacity or specialization. Our Curriculum-Guided Layer Scaling approach unifies prior ideas of curriculum learning and progressive model scaling into a single coordinated pretraining strategy.
2 Method
Figure 1: Curriculum-Guided Layer Scaling (CGLS) is a new paradigm for compute-efficient language model pretraining that grows data complexity and model depth in tandem. We illustrate CGLS for a Llama-3.2-1B scale model with four training stages. Training begins with an 8-layer model on a data split consisting equally of data from all levels (high-school, undergraduate, and graduate). The learned weights from this stage are transferred to a larger 10-layer model, freezing the pretrained weights and training the new layers on a small, balanced data split for better initialization. The entire model is then unfrozen and pretrained on the more difficult data split. This process is repeated until the target model scale is reached.
Humans do not learn language by memorizing the full dictionary on day one. Instead, we begin with simple patterns: repetitive sounds, basic grammar, and simple visual inputs, and gradually expand to handle more abstract concepts, longer sentences, and complex reasoning as our cognitive capacities mature. This progression in data complexity is tightly coupled with the brain’s developmental growth: as new neural structures form, they scaffold increasingly sophisticated representations of the world [29]. We study whether this notion of growing neural and data complexity in tandem can improve the compute efficiency of language model pretraining.
Previous work has explored increasing either neural or data complexity alone during language model pretraining, through curriculum learning and various approaches for expanding neural nets; we review these approaches in the Related Work above. Inspired by the structured development of the human brain, we develop and study a pretraining paradigm that synchronizes a data curriculum with the gradual addition of randomly initialized layers to the language model. We next describe these two components in our framework—curriculum and layer expansion—in turn.
Curriculum Learning.
To construct a pretraining curriculum, we organize documents into progressively harder splits, beginning with simpler and more structured samples before introducing more complex, diverse, and noisier data. Let $\mathcal{D}_{i}$ represent the dataset used at stage $i$, where $|\mathcal{D}_{i}|$ is the number of tokens. We first segment our pretraining dataset into easy, medium, and hard components $\mathcal{D}_{\text{Easy}},\mathcal{D}_{\text{Medium}}$ and $\mathcal{D}_{\text{Hard}}$. Then, the data distribution for stage $i$ can be expressed as a mixture of the easy, medium, and hard components with mixture weights $(p_{i},q_{i},r_{i}),p_{i}+q_{i}+r_{i}=1$:
$\mathcal{D}_{i}=p_{i}\cdot\mathcal{D}_{\text{Easy}}+q_{i}\cdot\mathcal{D}_{\text{Medium}}+r_{i}\cdot\mathcal{D}_{\text{Hard}}.$
The mixture weights $(p_{i},q_{i},r_{i})$ are adjusted to balance difficulty at each stage, as shown in Figure 1.
Layer Expansion.
In parallel with a structured curriculum, we expand the model’s capacity at each stage by incrementally adding transformer layers. Beginning with a shallow model (e.g., 6 layers), we train on simpler datasets before transferring the learned weights to a larger model with additional layers. Let $\mathbf{\Theta}^{(i)}$ represent the parameters of the model at stage $i$, where $\mathbf{\Theta}^{(i)}_{1:k}$ corresponds to the parameters of the first $k$ non-embedding layers. Then, at stage $i+1$, the parameters are initialized as:
$\mathbf{\Theta}^{(i+1)}:=\{\Theta^{(i)}_{\text{embed}}\>\circ\>\mathbf{\Theta}_{1:k}^{(i)}\>\circ\>\mathbf{\Theta}_{k+1:n}^{\text{random}}\>\circ\>\Theta^{(i)}_{\text{norm}}\>\circ\>\Theta^{(i)}_{\text{lm}}\}$
where $\mathbf{\Theta}_{k+1:n}^{\text{random}}$ represents the newly added, randomly initialized layers. Following the results of Kumar et al. [30], training is then performed in two phases:
  1. 1.
    Initialization phase: the new parameters $\mathbf{\Theta}_{k+1:n}^{\text{random}}$ are trained, while all other layers are frozen; that is, the embedding layer $\Theta^{(i)}_{\text{embed}}$, previous non-embedding layers $\mathbf{\Theta}^{(i)}_{1:k}$, normalization layer $\Theta^{(i)}_{\text{norm}}$, and language modeling head $\Theta^{(i)}_{\text{lm}}$ are all kept frozen.
  2. 2.
    Full tuning phase: all parameters in $\mathbf{\Theta}^{(i+1)}$ are optimized.
This approach aims to ensure that the representations learned in prior layers are effectively utilized and not overwritten when the model size is increased. This progression is also illustrated in Figure 1.
3 Experiments
The core intuition of our approach for compute-efficient pretraining is to follow a slow transition from highly structured to more diverse data while scaling model capacity. We construct pretraining curricula using the TinyStories [10], BookCorpus [53], and DataComp for Language Models (DCLM) [33] corpora. We first implement our approach at a GPT-2 Small scale [37], starting with a small model trained on simple, synthetic language patterns (TinyStories) and gradually increasing both data complexity and model size. Subsequently, we scale our approach to the parameter count of Llama-3.2-1B [9], applying it to a stratified version of the DCLM dataset, where document complexity levels (e.g., high school, undergraduate, graduate) are inferred via a DistilBERT classifier trained on a small set of GPT-4o-labeled samples.
Next, we describe this particular instantiation of data curricula that we used in our experiments.
3.1 Data Curricula
In conventional LLM pretraining, noisy and diverse datasets (e.g., Common Crawl or The Pile) are used early, and high-quality curated datasets are reserved for the final stages, or fine-tuning. This approach assumes that early exposure to broad linguistic patterns enhances generalization, while later fine-tuning reduces noise artifacts [6]. In contrast, we explore whether incorporating more structure in early stages and later transitioning to diverse corpora can improve downstream performance. We find that different curricula work best at our two different scales and describe them in depth next.
GPT-2-Small Scale Curriculum.
For the GPT-2 experiments, we start with a split containing samples primarily from TinyStories, consisting of synthetically-generated short stories comprehensible to a young child. The goal of training on such highly-structured synthetic data is to establish robust low-level linguistic representations. We then gradually introduce more challenging datasets, such as BookCorpus and DCLM, characterized by broader vocabulary, complex sentence structures, and diverse contexts. Our hypothesis is that early exposure to clean, consistent patterns helps the model develop reliable embeddings, which later support generalization to noisier data such as DCLM.
Llama-3.2-1B Scale Curriculum.
At the 1B scale, rather than using separate datasets of increasing complexity, we stratified the DCLM-Baseline-1.0 corpus itself into tiers of general, domain-specific, and specialized reasoning complexity. This enabled us to perform curriculum progression entirely within a single large corpus, while leveraging the model’s increased capacity to handle more complexity earlier in training as compared to the small-scale experiments. To implement this, we first perform complexity stratification using a separate pretrained LLM:
  1. 1.
    Constructing Classifier Training Set: Using GPT-4o-mini, we label a subset of 20,000 documents by classifying complexity into three levels:
    • High School: Structured text with limited abstraction or formal reasoning, such as news articles, beginner-level Wikipedia pages, or simple conversational stories.
    • Undergraduate: Content with moderate reading difficulty and some domain-specific concepts, including ArXiv abstracts and science articles that require basic abstract reasoning.
    • Graduate/Advanced: Highly technical or specialized content such as academic papers in AI/ML, legal documents, medical research studies, and code repositories (e.g., GitHub).
  2. 2.
    Class Balancing & Splitting: We then upsampled any underrepresented categories, and performed a random 80/10/10 split for training, validation, and testing, respectively.
  3. 3.
    Model-Based Classification: Finally, we trained a simple DistilBERT-based classifier [38] on this data, which achieved $>90\%$ test-set accuracy and was used to label the full 2M document subset of DCLM used for training.
GPT-2-Small (124M Parameters)
Stage CGLS (Simple) CGLS ($N_{1}=3$) CGLS ($N_{1}=6$)
Stage 1 100 / 0 / 0 70 / 10 / 20 70 / 10 / 20
Stage 2 0 / 100 / 0 25 / 25 / 50 15 / 25 / 60
Stage 3 0 / 0 / 100 20 / 20 / 60 5 / 5 / 90
Stage 4 5 / 5 / 90
New Layer Tuning 45 / 30 / 25 45 / 30 / 25
Table 1: Data curricula per training stage for the three GPT-2-Small parameter scale setups: CGLS (Simple), CGLS (starting model depth $N_{1}=3$), and CGLS ($N_{1}=6$). Each split is represented as a proportion of TinyStories, BookCorpus, and DCLM tokens with format TinyStories / BookCorpus / DCLM, used for tuning only the new layers (New Layer Tuning) or for full tuning of the entire model at that stage. The Simple approach trains on one dataset for each stage, for both the tuning of only the new-layers and for fine-tuning, while the $N_{1}=3$ and $N_{1}=6$ approaches combine datasets with mixture weights that vary across stages.
Llama-3.2-1B (1.2B Parameters)
Stage General (HS) Domain (UG) Specialized (Grad.)
Stage 1 25% 50% 25%
Stage 2 20% 40% 40%
Stage 3 15% 35% 50%
Stage 4 10% 10% 80%
New Layer Tuning 34% 33% 33%
Table 2: Data curricula per training stage for the Llama-3.2-1B parameter scale setups. The same set of optimized mixture weights are used for all CGLS experiments. Each stage is defined by mixture weights over three levels of complexity: high school–level (general understanding), undergraduate-level (some domain-specific knowledge), or graduate-level (highly specialized information).
Then, by choosing mixture weights over the stratified splits of DCLM, our data schedule (see Table 2) reflects a gradual shift from general to specialized content as model depth increases.
3.2 Compute-Matched Method Comparisons
In our experiments at the GPT-2-Small and LLaMA-3.2-1B parameter counts, we fix a compute budget in pretraining FLOPs across all methods to enable clean comparisons. While prior work on LLM scaling suggests that optimal training lies near the “Chinchilla” compute/data tradeoff curve [23], we select a lower FLOP budget due to academic compute constraints, using 700M tokens for the GPT-2-small scale experiments and 2.5B tokens for the Llama-3.2-1B scale experiments. Following well-established practices in the LLM scaling laws literature [26], we approximate the total compute budget for each experiment, in FLOPs, as:
[…]
We next detail the compute-matched configurations of CGLS that we evaluate, before discussing a set of natural baseline approaches applying curricula over data and model complexity.
CGLS Compute-Matched Configurations.
Our experiments consisted of training several compute-matched models under our CGLS framework, with varying dataset splits and starting number of layers $N_{1}$. These are summarized as follows, with detailed data splits shown in Tables 1 and 2.
  • Simple (GPT-2-Small scale only): Starting from a randomly-initialized 6 layer model, we train for one epoch on  [^1] M tokens from TinyStories. These weights are transferred to a 9 layer model, which is trained twice on 140M tokens of BookCorpus: once in the initialization phase (transferred layers frozen) and once in the full tuning phase (all layers unfrozen). We apply a similar progression at 12 layers, using 200M DCLM tokens for the initialization phase and 461M tokens for the full tuning phase.
  • CGLS ($N_{1}=C)$: These configurations follow a shared methodology across both model scales. Training begins with $C$ layers trained on a curriculum-balanced dataset, followed by a staged progression to the full model size. At each stage, newly added layers are first trained with a fixed data split (used for all initialization phases), before the full model is fine-tuned on a progressively more challenging data distribution.
Compute-Matched Baselines.
For both model scales, we evaluate the following compute-matched baselines under the same pretraining FLOP budget:
  • Randomized baseline: A model with the final depth (a “full model”) is trained from scratch on a randomized mixture of the available data.
  • Curricularized baseline: This baseline simulates a coarse curriculum by training the full model sequentially on increasingly complex data subsets. For GPT-2-Small, we train on TinyStories, then BookCorpus, then DCLM. For LLaMA-3.2-1B, we follow the same staged training but used the stratified DCLM splits (high school, undergraduate, graduate).
  • Layer-scaling only: This baseline incrementally expands the model depth across stages, like CGLS, but trains on the full (unstructured) data distribution at each stage without a curriculum. This isolates the contribution of progressive stacking alone.
  • MIDAS: MIDAS progressively stacks layers across training stages but does not control the complexity of the training data. We reimplemented the method from pseudocode provided in the original paper [39], as there is no public codebase, with PROP-3 stage budgets, a block size of 3 for GPT-2-Small, and a block size of 4 for LLaMA-3.2-1B.
Further CGLS Implementation Details.
At GPT-2-Small parameter scale, we scaled from 6 to 12 layers in 3 stages; at Llama-3.2-1B parameter scale, from 8 to 16 layers in 4 stages. The train FLOP budgets for each stage were optimized and allocated to favor later, larger stages (20% each for the first three stages; 40% for the final stage at Llama-3.2-1B scale), ensuring sufficient resources for fine-tuning on the most complex data. Additionally, we allocated 20% of the FLOPs for each stage to training only the new layers, and 80% for fine-tuning the full-unfrozen model. Additional details about our experimental setup can be found in A.3.
3.3 Evaluation
To measure generalization and reasoning capabilities, we evaluated the models on several downstream tasks:
  • PIQA (Physical Interaction QA) [5]: This dataset tests physical commonsense reasoning by presenting multiple-choice questions about everyday scenarios.
  • ARC Easy and Challenge [8]: The AI2 Reasoning Challenge benchmarks measure a model’s ability to answer elementary and middle-school science questions. The "Easy" subset contains relatively straightforward questions, while the "Challenge" subset consists of harder, multi-step reasoning tasks.
  • HellaSwag [50]: This task evaluates common-sense reasoning and next-word prediction, requiring the model to identify the most plausible continuation for a given context.
  • MMLU (Massive Multitask Language Understanding) [22]: MMLU consists of multiple-choice questions across 57 tasks spanning the humanities, STEM, social sciences, and more. Given our constrained academic budget for compute, raw accuracy on MMLU and other multiple-choice tasks may not reflect a meaningful signal. Instead, we evaluate Pointwise Mutual Information (PMI) Accuracy and Token-Normalized Accuracy [32, 19], metrics that better account for length and bias effects in generative modeling to obtain a low-variance signal at small scales. These metrics are defined as:
    $\text{acc}_{\text{pmi}}=\arg\max_{i}\ln\frac{P(a_{i}\mid q)}{P(a_{i}\mid u)},\quad\text{acc}_{\text{token}}=\arg\max_{i}\frac{\ln P(a_{i}\mid q)}{\texttt{num\_tokens}(a_{i})},$
[…]
  • LAMBADA [36]: An open-book language modeling task designed to assess a model’s ability to perform long-range dependency reasoning. We report zero-shot accuracy, computed as whether the model correctly predicts the final word of a passage.
All tasks are evaluated in a zero-shot setting to measure base model quality without additional fine-tuning. We report the token-normalized accuracy, as described above, for all multiple-choice tasks, and PMI for MMLU. We additionally compute perplexity on a 180,000 sample subset of the Pile dataset, capturing how well the methods model diverse, noisy data from a variety of domains.
4 Results
Our experiments evaluate the performance of the CGLS training strategy at two parameter counts (matched to GPT-2-Small and Llama-3.2-1B) across perplexity and zero-shot accuracy on several downstream tasks. Table 3 summarizes the findings, comparing CGLS to several baselines with matched compute budgets in FLOPs. Across both model scales, CGLS consistently outperforms baseline approaches on reasoning-focused benchmarks such as ARC and PIQA, demonstrating its effectiveness in improving generalization under fixed compute budgets. CGLS shows particular strength in structured and knowledge-intensive tasks, supporting the benefits of aligning model growth with data complexity.
Model
Pile
Perplexity
PIQA ARC-E ARC-C HellaSwag Lambada
MMLU
Average
GPT-2-Small
Pretrained GPT-2-Small 23.97 62.51% 39.48% 22.70% 31.14% 38.96%
Baseline (Randomized) 37.51 58.00% 36.11% 21.08% 26.60% 35.45%
Baseline (Curricularized) 36.09 59.09% 36.78% 22.18% 26.43% 36.12%
Layer Scaling Only 39.67 58.76% 35.82% 21.42% 26.35% 35.59%
MIDAS [39] 49.90 56.91% 34.81% 21.16% 25.66% 34.64%
CGLS (Simple) 35.93 59.03% 37.67% 21.76% 19.50% 34.49%
CGLS ($N_{1}=3$) 35.48 59.19% 37.46% 21.08% 26.77% 36.13%
CGLS ($N_{1}=6$) 34.33 60.12% 36.62% 22.53% 26.55% 36.46%
Llama-3.2-1B
Pretrained Llama-3.2-1B 9.25 74.32% 60.56% 36.52% 63.67% 62.00% 31.50% 54.76%
Baseline (Randomized) 20.83 59.09% 36.36% 24.24% 34.20% 32.99% 22.11% 34.80%
Baseline (Curricularized) 21.93 58.27% 35.19% 25.43% 34.65% 33.09% 24.63% 35.25%
Layer Scaling Only 29.53 57.45% 32.58% 23.46% 30.00% 27.73% 22.62% 32.56%
MIDAS [39] 27.64 55.44% 32.58% 24.23% 29.62% 25.87% 23.81% 31.97%
CGLS ($N_{1}=6$) 26.58 61.70% 40.03% 23.55% 34.20% 28.92% 23.48% 36.28%
CGLS ($N_{1}=8$) 25.34 61.21% 41.37% 25.43% 34.65% 32.74% 26.55% 36.97%
Table 3: Evaluation results comparing CGLS and baseline approaches across average perplexity on The Pile dataset and core reasoning- and knowledge-intensive downstream tasks. CGLS, particularly the configuration starting with half of the final model depth, improves accuracy on downstream tasks like PIQA compared to the baselines. All models are compute-matched, but utilize much less compute than the pretrained variants of these models. Not compute-matched to the other methods, and trained on orders-of-magnitude more data. We provide pretrained results to illustrate a rough performance upper bound for the given architecture. The best model is bolded for each benchmark, with the second-best model underlined.
GPT-2-Small Scale.
The GPT-2 scale CGLS models achieve consistent improvements on downstream tasks (Table 3), and consistently lower perplexity across nearly all subsets of The Pile, as illustrated in Figure 2. For example, the $N_{1}=6$ model attains a 60.12% accuracy on PIQA, a 2.1-point improvement over the randomized baseline, along with gains on ARC-Challenge (22.53% vs. 21.08%).
LLaMA-3.2-1B Scale.
The larger models demonstrate clearer trends. While perplexity results diverged from the GPT-2-Small trend, with the randomized baseline achieving the lowest perplexity across all Pile domains (see A.1 for detailed perplexity results), CGLS nonetheless achieves the strongest downstream performance. CGLS with $N_{1}=8$ achieves the highest accuracy across nearly all downstream tasks, improving over the best baseline on PIQA (61.21% vs. 59.09%), ARC-Easy (41.37% vs. 36.36%), and MMLU-STEM (26.44% vs. 24.84%). The gains on ARC-Easy were especially striking, exceeding the best baseline by over 5 percent. Notably, the stacking baselines, without a synchronized data curricula, did not improve downstream performance. For MMLU, results remained near chance overall, reflecting the well-established difficulty of the benchmark for small models [32, 19]. However, token-normalized and PMI metrics revealed some structure. The STEM category, in particular, showed a meaningful increase: 26.44% for CGLS vs. 24.84% for the curricularized baseline and 21.92% for the randomized baseline. This suggests that CGLS may be particularly helpful in technical and reasoning-heavy domains. Further analysis is needed to understand rankings within MMLU subsets (see A.2 for further analysis), but the trend is promising.
Figure 2: Model perplexity at GPT-2-Small parameter count across various individual subsets of The Pile, for three CGLS configurations and the strongest two baselines. Lower perplexity scores indicate better performance. CGLS consistently outperforms the baselines, particularly on complex knowledge-intensive datasets such as ArXiv, PubMed, and NIH.
Comparing Progressive Scaling Strategies.
Among the GPT-2 CGLS variants, the best performance was consistently achieved with the $N_{1}=6$ model, starting from half the final model depth. This configuration outperformed both the ’Simple’ strategy and the $N_{1}=3$ model, suggesting that initializing from too small a model may hinder the transferability of learned representations. Similarly, for LLaMA-3.2-1B, CGLS with $N_{1}=8$ (half of the final 16-layer model) delivered the best downstream performance. Increasing the starting depth beyond half the final depth did not yield additional improvements (see A.1). This suggests that starting from half the model depth may strike a near-optimal balance between early-stage tractability and representational richness in progressive training frameworks.
5 Discussion
Inspired by how humans acquire cognitive capabilities in tandem with brain maturation and exposure to progressively more complex information, we align model capacity with data complexity during language model pretraining. We propose a new pretraining paradigm, Curriculum-Guided Layer Scaling (CGLS), that grows models in depth while adapting the distribution of training data to match their evolving capabilities. Applied to pretraining at GPT-2-Small and LLaMA-3.2-1B parameter counts, our approach yields consistent improvements in downstream task performance. In compute-matched comparisons, CGLS models outperform baselines on benchmarks requiring structured reasoning and domain-specific understanding, such as ARC, PIQA, and MMLU-STEM. Importantly, stacking-based baselines that expand model depth without synchronized data curricula do not improve downstream performance—CGLS significantly outperforms them across key benchmarks, demonstrating that aligning model growth with data complexity can unlock the benefits of progressive growing. Altogether, our results establish CGLS as a promising new paradigm for compute-efficient language model pretraining.
Limitations.
While CGLS yields improvements across several knowledge-intensive and reasoning tasks, in a few tasks such as LAMBADA and HellaSwag, CGLS models perform comparably to the baselines, suggesting that not all tasks benefit equally from the progressive curriculum. We also observe that CGLS does not always lower perplexity: at LLaMA-3.2-1B scale, the randomized baseline achieved slightly better perplexity despite underperforming on reasoning tasks. This disconnect between perplexity and functional generalization highlights the need for better evaluation frameworks.
Future Work.
We find in experiments that the initial model size is a crucial hyperparameter: starting with half the final number of layers (“$N_{1}=N/2$”) yielded the best performance at both pretraining scales. Starting too small (e.g., $N_{1}=3$ at GPT-2 scale or $N_{1}=6$ at Llama-3.2-1B scale) led to weaker final models, which we postulate may stem from insufficient capacity to master basic linguistic elements. Starting too large (e.g., $N_{1}=10$ for LLaMA) offered no additional benefit and, in some cases, underperformed standard training (results provided in A.1). Given our limited academic budget, we were not able to exhaustively tune these and other hyperparameters, suggesting that CGLS may be further improved. Future work could explore several directions: (1) more finely tuning the stage lengths and dataset proportions, (2) exploring other notions—e.g., syntactic diversity, lexical richness, or factuality—for defining the curriculum, (3) applying curriculum-aware optimization schedules (e.g., varying learning rates with data complexity), and (4) applying CGLS to non-text modalities (e.g., vision or multimodal data).
Broader Impacts.
CGLS is a general-purpose pretraining framework designed to improve the compute efficiency of language model pretraining. Potential positive impacts include democratizing access to high-performing language models, reducing pretraining compute for research groups with limited resources, and improving model interpretability through more structured growth. However, like other LLM training methods, CGLS could also be used to accelerate development of models deployed in harmful or unregulated ways. These include generating disinformation, bias amplification, or surveillance applications. While CGLS itself does not introduce new capabilities in these domains, it could reduce the cost and time needed to train capable models, slightly lowering the barrier for misuse. To mitigate risks, we recommend applying existing safety and alignment techniques during downstream fine-tuning stages, particularly for open-domain deployment. We also stress that our approach is most beneficial when paired with curated datasets and clear usage constraints.
Appendix A Technical Appendices and Supplementary Material
A.1 Additional Results
We present a more detailed comparison, including results for CGLS starting at 10 layers, in Table 4. CGLS with $N_{1}=8$ achieves the best overall downstream performance, including the highest accuracies on PIQA, ARC-E, ARC-Challenge, Hellaswag, and MMLU-STEM. CGLS with $N_{1}=6$ performs competitively, whilst starting with too large a model ($N_{1}=10$) underperforms on both perplexity and accuracy, supporting our hypothesis that synchronizing capacity with complexity is crucial.
Figure 3 further illustrates this trend, showing per-domain perplexity across The Pile. The randomized baseline achieves the lowest perplexity across all subsets, but CGLS outperforms other baselines such as MIDAS and layer-scaling without curricula, particularly on reasoning-heavy or technical subsets like Github and StackExchange.
Model
Pile
Perplexity
PIQA ARC-E ARC-C HellaSwag Lambada
MMLU
Llama-3.2-1B
Baseline (Randomized) 20.83 59.09% 36.36% 24.24% 34.20% 32.99% 22.11% 34.80%
CGLS ($N_{1}=6$) 26.58 61.70% 40.03% 23.55% 34.20% 28.92% 23.48% 36.28%
CGLS ($N_{1}=8$) 25.34 61.21% 41.37% 25.43% 34.65% 32.74% 26.55% 36.97%
CGLS ($N_{1}=10$) 31.69 57.34% 35.19% 23.38% 29.29% 24.96% 24.78% 32.49%
Table 4: Evaluation results comparing the compute-matched randomized baseline to CGLS across initial model sizes on The Pile perplexity and downstream tasks: PIQA, ARC-Easy, ARC-Challenge, Hellaswag, Lambada, and MMLU-STEM. The approach starting with half of the final model depth, achieves the highest accuracy among the CGLS experiments. The best model is bolded for each benchmark, with the second-best model underlined.
Figure 3: Model perplexity across various individual subsets of The Pile for all baselines versus CGLS for Llama-3.2-1B. Lower perplexity scores indicate better performance. Across all subsets, the randomized baseline achieves the lowest perplexity, with progressive scaling models outperforming MIDAS and layer scaling without a data curriculum.
Baseline
Randomized
Baseline
Curricularized
Layer
Scaling
MIDAS
CGLS
($N_{1}=6$)
CGLS
($N_{1}=8$)
Average 23.17% 24.93% 24.37% 23.50% 23.82% 25.10%
Other 24.35% 27.75% 27.18% 26.27% 27.63% 27.52%
Professional 24.20% 25.43% 25.66% 23.39% 25.38% 25.59%
Humanities 23.45% 24.72% 24.94% 24.56% 24.18% 23.99%
STEM 22.11% 24.63% 22.62% 23.81% 23.48% 26.55%
Social Sciences 21.98% 26.79% 23.90% 22.27% 23.07% 22.99%
biology 22.78% 22.78% 25.00% 22.92% 27.22% 27.78%
business 25.69% 25.21% 27.50% 22.92% 27.38% 26.30%
chemistry 18.81% 25.41% 22.44% 25.41% 20.46% 24.75%
computer science 26.21% 25.49% 23.54% 26.21% 23.06% 27.67%
economics 21.43% 27.09% 25.20% 20.89% 24.39% 25.07%
engineering 22.76% 22.76% 24.14% 25.52% 25.52% 26.21%
geography 18.69% 27.27% 18.69% 15.66% 18.69% 19.70%
health 23.05% 23.96% 26.89% 22.01% 26.22% 26.59%
history 23.44% 27.20% 24.73% 24.52% 23.87% 24.84%
law 24.56% 26.89% 23.65% 24.90% 23.65% 24.33%
math 20.49% 25.09% 20.11% 23.50% 21.43% 27.44%
other 24.35% 27.75% 27.18% 26.27% 27.63% 27.52%
philosophy 23.21% 23.71% 25.55% 24.65% 24.40% 23.56%
physics 22.81% 25.47% 23.28% 22.66% 23.91% 23.91%
politics 20.63% 25.46% 21.38% 21.75% 18.96% 19.52%
psychology 23.37% 27.10% 25.00% 24.30% 24.69% 23.76%
sociology 25.87% 23.38% 19.90% 23.88% 23.38% 24.38%
Table 5: MMLU accuracies across broader categories and finer subcategories for all Llama-3.2-1B scale models. We report zero-shot accuracy using PMI-based evaluation. The best-performing model for each category or subcategory is bolded, and the second-best is underlined. While the differences are often marginal, CGLS achieves the highest average accuracy and shows notable gains in STEM and reasoning-heavy subdomains.
A.2 MMLU Breakdown
Table 5 presents a finer-grained view of MMLU performance across individual categories and subcategories. While overall accuracy remains near chance, CGLS with $N_{1}=8$ achieves the highest average accuracy (25.10%), outperforming all baselines. The strongest gains appear in technical domains such as STEM (26.55%), math (27.44%), and computer science (27.67%), suggesting that progressive scaling may enhance performance in reasoning-heavy subfields. However, performance varies substantially across subcategories, with different models achieving the best result depending on the domain.
A.3 Experimental Setup
All training was performed using the Hugging Face Transformers library [46] on NVIDIA H100 80GB GPUs with NVLink. Each model was initialized with a learning rate of $2\times 10^{-4}$ and optimized using AdamW with a weight decay of 0.01. A warmup-stable-decay (WSD) learning rate schedule was applied, with 1,000 warmup steps during the initial stage (smallest model), followed by a stable learning rate across intermediate stages, and cosine decay to zero during the final stage (largest model). Layers added after model expansion were initially trained with a higher learning rate of $5\times 10^{-4}$, for 20% of the compute budget for the given stage. Optimizer states were reset between stages for all experiments involving layer expansion.
A.4 Dataset and Model Licenses
This work makes use of several publicly available datasets and pretrained models. The TinyStories dataset is released under the CDLA-Sharing-1.0 license and is composed of synthetic data generated for research purposes. The BookCorpus dataset, though widely used in academic NLP research, is no longer distributed and therefore does not have a formal license, though it may include copyrighted content. We include it here in line with standard usage in prior literature, but note its ambiguous legal status.
The DCLM (DataComp for Language Models) dataset is released under the Creative Commons Attribution 4.0 license as part of the DataComp benchmark suite, with a clear focus on research usability and content quality filtering.
For pretrained models, we use GPT-2-Small, which is released under the MIT model license. The LLaMA-3.2-1B model is made available by Meta under the LLaMA 3.2 Community License. While usable for research and certain commercial applications, the license prohibits redistribution of any model weights generated by Meta.